# Exact Thermal Eigenstates of Nonintegrable Spin Chains at Infinite Temperature

### v2

@@ Abstract
Discussion.— We have studied a class of volume-law states, which we call entangled antipodal pair (EAP) states, and thoroughly characterized these states by providing the necessary and sufficient conditions that Hamiltonians must satisfy in order to have an EAP state as an eigenstate. Moreover, we have rigorously shown that some of such Hamiltonians are nonintegrable. Our EAP states are indistinguishable from the Gibbs state at infinite temperature. In other words, we have written down, for the first time, analytic expressions for thermal energy eigenstates of nonintegrable many-body systems. We have also devised a method for constructing finite-temperature thermal states using an EAP state.
@@
Some readers may be concerned that the energy eigenspace containing the EAP state is excessively large because, with an exponentially large number of states, it can be not so difficult to obtain a highly entangled state as their linear combination, even when each is hardly entangled. However, numerical results show that in Model 1, the degeneracy of the EAP state is only two in the momentum sector  [^6] Furthermore, by extending the locality of interactions from two to three, we have also obtained an example of a pair of the EAP eigenstate and the nonintegrable Hamiltonian that does not have degeneracy in the corresponding momentum sector [27]. In addition, we can also confirm that all states in the degenerate eigenspace obey a volume law with the maximal coefficient $\log 2$. Hence, the nontriviality of our findings is not diminished by the degeneracy.
@@
Our result suggests that theoretical analysis may be feasible even for thermal eigenstates of nonintegrable systems and may pave the way for giving a provable example of the ETH. Since EAP states themselves are at infinite temperature, constructing finite temperature thermal eigenstates remains challenging. However, as we have shown, one can obtain thermal states at arbitrary temperature by applying an imaginary time evolution, a low-complexity operation that can be well approximated via a matrix product operator with small bond dimension [37], to the EAP state. This implies that the expressivity of states derived from EAP states is quite high. Thus, EAP states would be one of the most promising starting points for constructing thermal eigenstates at finite temperature. Furthermore, it will provide theoretical methodologies not only for thermalization, but also for various areas of quantum statistical mechanics that use thermal pure states, such as the formulation of statistical mechanics and finite temperature simulations.
@@
Acknowledgement.— We are grateful to N. Shiraishi, A. Shimizu, H. Katsura, K. Fujii, K. Mizuta, H. Kono, M. Yamaguchi, and Z. Wei for useful discussions. We would like to especially thank H. K. Park for careful reading and valuable comments on a draft of this paper. YY is supported by the Special Postdoctoral Researchers Program at RIKEN. YC is supperted by Japan Society for the Promotion of Science KAKENHI Grant No. JP21J14313.
Note added: After the completion of this work, we became aware of a recent related work by A. N. Ivanov and O. I. Motrunich (arXiv:2403.05515) which constructs an exact volume-law eigenstate of the PXP and related models.
@@ I Proof of Theorem 1
where $Y=X+N/2$ and $\nu_{k}=\mu_{k-N/2}$ for $k\in Y$. Since $\omega_{j}^{\mu_{j}}=\pm 1$, this implies Eq. (6).
@@
Remark: The above proof also shows that, if we only need to show (iii) $\Rightarrow$ (ii), we can relax the condition on the locality of interactions in the main text, “$J_{X}^{\vec{\mu}}=0$ for any subset $X$ with $D(X)\geq N/4$” to “$J_{X}^{\vec{\mu}}=0$ for any subset $X$ with $D(X)\geq N/2$.”
II List of translation-invariant and nearest-neighbor-interacting Hamiltonians
We provide a list of translation-invariant (for single-site translations) and nearest-neighbor-interacting Hamiltonians having an EAP state as an eigenstate. First, we exclude the case of free spins, as it is trivial. Next, for some cases where only one of $J^{\mu\nu}$ is non-zero, we find that $2^{N/2}$ EAP states are degenerate, so we also exclude such cases. Then pairs of the EAP state and the Hamiltonian are limited to the following five types (and their equivalents obtained by appropriate permutations of directions of the Pauli matrices). It can be readily confirmed through direct calculations that the pairs of the EAP state and the Hamiltonian listed below satisfy Eq. (6).
@@ II.2 Case where the EAP state is invariant under $2$-site translation
The EAP state |$\ket{2;10,01}$ is an eigenstate of the Hamiltonian defined by
@@ II.3 Case where the EAP state is invariant under $3$-site translation
The EAP state |$\ket{3;10,11,01}$ is an eigenstate of the Hamiltonian defined by
@@ II.4 Case where the EAP state is invariant under $4$-site translation
The EAP state |$\ket{4;10,11,00,01}$ is an eigenstate of the Hamiltonian defined by
@@ III Proof of Theorem 3
This means that $\hat{Q}$ is a $(k-1)$-local conserved quantity. Applying the same argument to $k-1$, $k-2$,…, and $3$-local conserved quantity, we can show that any $k$-local conserved quantity with $k\leq N$ have to be a $2$-local conserved quantity.
@@
This means that $\hat{Q}$ is a $(k-1)$-local conserved quantity. Applying the same argument to $k-1$, $k-2$,…, and $3$-local conserved quantity, we can show that any $k$-local conserved quantity with $k\leq N$ have to be a $2$-local conserved quantity.
@@ VI.2 Nondegenerate Hamiltonian with next-nearest-neighbor interactions
We also verify the nonintegrability of model (S221). Because of the existence of $J^{zzz}$, model (S221) is not mapped to a free fermionic (integrable) system by the Jordan-Wigner transformation. We confirm nonintegrability of model (S221) by calculating the distribution of the ratio of consecutive level spacings, as in Fig. S1. Figure S4 plots the distribution constructed from energy eigenvalues in the eigenspace of translation $\hat{\mathcal{T}}$ and $\mathbb{Z}_{2}$ transformation $\bigotimes_{j=1}^{N}\hat{\sigma}_{j}^{z}$ with momentum $k=\pi/2$ and parity $\mathcal{P}=\pm 1$. This plot is well described by the Gaussian unitary ensemble distribution [45] (dashed line) and well separated from the Poisson distribution [45] (dotted line), indicating the nonintegrability of the model.

### v4

@@ Abstract
Discussion— We have studied a class of volume-law states, which we call entangled antipodal pair (EAP) states, and thoroughly characterized these states by providing the necessary and sufficient conditions that Hamiltonians must satisfy in order to have an EAP state as an eigenstate. Moreover, we have rigorously shown that some of such Hamiltonians are nonintegrable. Our EAP states are indistinguishable from the Gibbs state at infinite temperature. In other words, we have written down analytic expressions for thermal energy eigenstates of nonintegrable many-body systems. We have also devised a method for constructing finite-temperature thermal states using an EAP state.
@@
Some readers may be concerned that the energy eigenspace containing the EAP eigenstate is excessively large because, with an exponentially large number of states, it can be not so difficult to obtain a highly entangled state as their linear combination, even when each is hardly entangled [49]. However, additional numerical calculations [35] show that in Model 1, the degeneracy at zero energy is only $2$ in the momentum sector  [^8] Furthermore, by extending the locality of interactions from two to three, we have also obtained an example of a pair of the EAP eigenstate and the nonintegrable Hamiltonian that does not have degeneracy in the corresponding momentum sector [35].. In addition, we can also confirm that all states in the degenerate eigenspace have a volume-law entanglement with the maximal coefficient $\log 2$ [35]. Hence, the nontriviality of our findings is not diminished by the degeneracy.
@@
Our result suggests that theoretical analysis may be feasible even for thermal eigenstates of nonintegrable systems and may pave the way for giving a provable example of the ETH. Since EAP states themselves can give only a few of eigenstates at infinite temperature, constructing general eigenstates remains challenging. However, as we have shown, one can obtain thermal states at arbitrary temperature by applying an imaginary-time evolution, a low-complexity operation that can be well approximated via a matrix product operator with small bond dimension [51], to the EAP state. This implies that the expressivity of states derived from EAP states is quite high. Thus, EAP states would be one of the most promising starting points for constructing general eigenstates, including those at finite temperature. Furthermore, it will provide theoretical methodologies not only for thermalization, but also for various areas of quantum statistical mechanics that use thermal pure states, such as the formulation of statistical mechanics and finite temperature simulations.
@@
Acknowledgments— We are grateful to N. Shiraishi, A. Shimizu, H. Katsura, K. Fujii, K. Mizuta, H. Kono, M. Yamaguchi, and Z. Wei for useful discussions. We would like to especially thank H. K. Park and A. N. Ivanov for careful reading and valuable comments on a draft of this Letter. Y.Y. is supported by the Special Postdoctoral Researchers Program at RIKEN. Y.C. is supported by Japan Society for the Promotion of Science KAKENHI Grant No. JP21J14313, the Special Postdoctoral Researchers Program at RIKEN, and JST ERATO Grant No. JPMJER2302, Japan.
@@ I Proof of Theorem 1
where $Y=X+N/2$ and $\nu_{k}=\mu_{k-N/2}$ for $k\in Y$. Since $(\omega_{j}^{\mu_{j}})^{-1}=\omega_{j}^{\mu_{j}}$, this implies Eq. (7).
@@
Remark: The above proof also shows that, if we only need to show (iii) $\Rightarrow$ (ii), we can relax the condition on the locality of interactions in the main text, “$J_{X}^{\vec{\mu}}=0$ for all subset $X$ with $D(X)\geq N/4$” to “$J_{X}^{\vec{\mu}}=0$ for all subset $X$ with $D(X)\geq N/2$.”
II List of translation-invariant and nearest-neighbor-interacting Hamiltonians (Theorem 2)
We provide a list of translation-invariant (for single-site translations) and nearest-neighbor-interacting Hamiltonians having an EAP state as an eigenstate. First, we exclude the case of free spins, as it is trivial. Next, for some cases where only one of $J^{\mu\nu}$ is non-zero, we find that $2^{N/2}$ EAP states are degenerate, so we also exclude such cases. Then pairs of the EAP state and the Hamiltonian are limited to the following five types (and their equivalents obtained by appropriate permutations of directions of the Pauli matrices). It can be readily confirmed through direct calculations that the pairs of the EAP state and the Hamiltonian listed below satisfy Eq. (7) or (8).
@@ II.2 Case where the EAP state is invariant under $2$-site translation
The EAP state |$\ket{2;01,10}$ is an eigenstate of the Hamiltonian defined by
@@ II.3 Case where the EAP state is invariant under $3$-site translation
The EAP state |$\ket{3;01,10,11}$ is an eigenstate of the Hamiltonian defined by
@@ II.4 Case where the EAP state is invariant under $4$-site translation
The EAP state |$\ket{4;00,01,10,11}$ is an eigenstate of the Hamiltonian defined by
@@ III Proof of Theorem 3
This means that $\hat{Q}$ is a $(k-1)$-local conserved quantity. Applying the same argument to $k-1$, $k-2$,…, and $3$-local conserved quantity, we can show that any $k$-local conserved quantity with $k\leq N/2$ have to be a $2$-local conserved quantity.
@@
This means that $\hat{Q}$ is a $(k-1)$-local conserved quantity. Applying the same argument to $k-1$, $k-2$,…, and $3$-local conserved quantity, we can show that any $k$-local conserved quantity with $k\leq N/2$ have to be a $2$-local conserved quantity.
@@ VI.2 Nondegenerate Hamiltonian with next-nearest-neighbor interactions
We also verify the nonintegrability of model (S221). Because of the existence of $J^{zzz}$, model (S221) is not mapped to a free fermionic (integrable) system by the Jordan-Wigner transformation. We confirm nonintegrability of model (S221) by calculating the distribution of the ratio of consecutive level spacings, as in Fig. S1. Figure S4 plots the distribution constructed from energy eigenvalues with eigenmomentum $k=\pi/2$ and parity $\mathcal{P}=\pm 1$. (Here the eigenmomentum $k$ is related to the eigenvalue of translation $\lambda$ as $\lambda=e^{-ik}$, and the parity $\mathcal{P}$ is the eigenvalue of $\mathbb{Z}_{2}$ symmetry $\bigotimes_{j=1}^{N}\hat{\sigma}_{j}^{z}$.) This plot is well described by the Gaussian unitary ensemble distribution [59] (dashed line) and well separated from the Poisson distribution [59] (dotted line), indicating the nonintegrability of the model.
@@
VII Relation to previously constructed thermal eigenstates
In this section, we discuss the relation between our Theorem 1 and previously constructed thermal eigenstates of nonintegrable models [36, 37].
VII.1 Relation to the results by A. Udupa, S. Sur, et al.
Now we discuss relation to the results by A. Udupa, S. Sur, et al. [36]. They investigated the model described by the Hamiltonian
$\hat{H}_{ZZZ}=\sum_{j=1}^{N}\bigl{(}J\hat{\sigma}_{j}^{z}\hat{\sigma}_{j+1}^{z}\hat{\sigma}_{j+2}^{z}+h\hat{\sigma}_{j}^{x}\bigr{)}$ (S224)
with the periodic boundary condition. This model is expected to be nonintegrable because its level spacing distribution is described by the Wigner-Dyson distribution [36].
They showed that $\hat{H}_{ZZZ}$ has an EAP state |$\ket{1;11}$ as an eigenstate with the eigenvalue 00. This fact can be readily verified by checking Eq. (7) and using (iii) $\Rightarrow$ (ii) of our Theorem 1.
Note that $\hat{H}_{ZZZ}$ has exponentially large degeneracy, which is greater than or equal to $2^{N/2}$ as they showed, at the eigenvalue 00. They constructed some degenerate eigenstates other than |$\ket{1;11}$, but not all of them. Thus, their results do not exclude the possibility that, for another basis of zero energy eigenspace, each of the basis vectors becomes hardly entangled. (By contrast, our Fig. S3 shows that any superpositions of degenerate states have volume law entanglement with coefficients almost the same as the maximal value $\log 2$.)
Note also that, although they argued that constructed eigenstates including |$\ket{1;11}$ are quantum many-body scar states, the EAP state |$\ket{1;11}$ is thermal in the most fundamental sense as explained in the main text. (We agree that some other degenerate eigenstates that contain singlet pairs between neighboring sites [36] are many-body scar states because they can be distinguished from the Gibbs state at $\beta=0$ by a local observable on the neighboring sites.)
VII.2 Relation to the results by A. N. Ivanov and O. I. Motrunich
Next, we discuss the relation to the results by A. N. Ivanov and O. I. Motrunich [37]. They investigated the PXP model, which is described by the Hamiltonian
$\hat{H}_{PXP}=\sum_{j=1}^{N}\hat{P}_{j-1}\hat{\sigma}_{j}^{x}\hat{P}_{j+1}$ (S225)
with the periodic boundary condition, where $\hat{P}_{j}=(1+\hat{\sigma}_{j}^{z})/2$. Let $\hat{\mathcal{P}}_{\mathrm{Ryd}}$ be the projection operator to the nearest neighbor Rydberg blockaded subspace, which is spanned by the computational basis vectors |$\ket{\vec{\sigma}}$ whose bitstrings $\vec{\sigma}$ do not contain neighboring “$11$.” [Here, the computational basis is defined by Eq. (S205).] This projection operator can be written as
$\hat{\mathcal{P}}_{\mathrm{Ryd}}=\prod_{j=1}^{N}(1-\hat{Q}_{j}\hat{Q}_{j+1}),$ (S226)
where $\hat{Q}_{j}=1-\hat{P}_{j}$ and the periodic boundary condition is imposed again.
They showed that the projection of an EAP state, $\hat{\mathcal{P}}_{\mathrm{Ryd}}\ket{1;01}$, is an eigenstate of $\hat{H}_{PXP}$ (up to normalization) with the eigenvalue 00. This fact can be verified from our Theorem 1 as follows. By checking Eq. (7) and using (iii) $\Rightarrow$ (ii) of our Theorem 1, we can show that the EAP state |$\ket{1;01}$ itself satisfies $\hat{H}_{PXP}\ket{1;01}=0$. Furthermore, we can show that each local term $\hat{P}_{j-1}\hat{\sigma}_{j}^{x}\hat{P}_{j+1}$ in $\hat{H}_{PXP}$ commutes with $\hat{\mathcal{P}}_{\mathrm{Ryd}}$ by naively calculating the commutator, and hence $\hat{H}_{PXP}$ also commutes with $\hat{\mathcal{P}}_{\mathrm{Ryd}}$. Thus we have
$\hat{H}_{PXP}\hat{\mathcal{P}}_{\mathrm{Ryd}}\ket{1;01}=\hat{\mathcal{P}}_{\mathrm{Ryd}}\hat{H}_{PXP}\ket{1;01}=0,$ (S227)
which corresponds to the result explained above.
In addition, we show in the following that the state $\hat{\mathcal{P}}_{\mathrm{Ryd}}\ket{1;01}$ is thermal in the sense that it is indistinguishable from the maximally mixed state of the Rydberg blockaded subspace, whose density matrix is given by $\hat{\mathcal{P}}_{\mathrm{Ryd}}$ up to normalization. We examine the reduced density matrix on consecutive $m$ sites $\{1,...,m\}$ ($m\leq N/2$) for both the state $\hat{\mathcal{P}}_{\mathrm{Ryd}}\ket{1;01}$ and the maximally mixed state, and compare them. Let $\mathcal{F}_{m}^{(\mathrm{o})}$ be the set of bitstrings that satisfy the Rydberg blockaded condition for the open boundary condition. As shown in Sec. II of Supplemental Material of Ref. [37], the reduced density matrix on sites $\{1,...,m\}$ ($m\leq N/2$) for the state $\hat{\mathcal{P}}_{\mathrm{Ryd}}\ket{1;01}$ is given by (up to normalization)
$\sum_{\vec{\sigma}\in\mathcal{F}_{m}^{(\mathrm{o})}}F_{\frac{N}{2}+2-m-\sigma_{1}-\sigma_{m}}\ket{\vec{\sigma}}\bra{\vec{\sigma}},$ (S228)
where $F_{n}$ is the $n$-th Fibonacci number. On the other hand, by a calculation similar to theirs, we can show that the reduced density matrix on the same sites for the maximally mixed state is given by (up to normalization)
$\sum_{\vec{\sigma}\in\mathcal{F}_{m}^{(\mathrm{o})}}F_{N+2-m-\sigma_{1}-\sigma_{m}}\ket{\vec{\sigma}}\bra{\vec{\sigma}}.$ | . (S229)
Hence, in the limit of $N\to\infty$ (with fixed $m$), Eqs. (S228) and (S229) give the same reduced density matrices. This indicates that $\hat{\mathcal{P}}_{\mathrm{Ryd}}\ket{1;01}$ is indistinguishable from the maximally mixed state by the expectation values of any local observables.
Thus, the state $\hat{\mathcal{P}}_{\mathrm{Ryd}}\ket{1;01}$ studied in Ref. [37] is also an example of exact thermal eigenstates of nonintegrable systems (because the PXP model is known to be nonintegrable [57]), although they referred to the state $\hat{\mathcal{P}}_{\mathrm{Ryd}}\ket{1;01}$ as a quantum many-body scar state. As explained in the main text, our terminology is based on one of the standard notions of thermal equilibrium, “MITE [3].”